# LibreLog: Accurate and Efficient Unsupervised Log Parsing Using Open-Source Large Language Models

### v2

@@ V Evaluation
During retrieval augmented log parsing, varying the number of selected logs affects the performance of OpenLogParser. Retrieving three logs into the prompt yields the highest effectiveness. Fig 4 shows the OpenLogParser performance with variations in the number of logs from a group retrieved into the prompts. OpenLogParser maintains high accuracy across various sample sizes, with optimal performance achieved when the sample size is set to three, reaching peak values in both GA and PA. Notably, when the sample size is reduced to one, GA and PA drop to 0.80 and 0.70, respectively, representing a decline of 8.26% and 18% compared to a sample size of three. This reduction highlights the challenges LLM faces in parsing logs accurately without sufficient comparative data, such as multiple log comparisons or labeled logs. As the sample size increases from one to two, both GA and PA show significant improvements, peaking when the sample size reaches three. However, further increases in sample size from three to eight result in slight decreases in GA and PA, stabilizing around 0.865 and 0.835, respectively. This suggests that an excess of log samples may introduce noise, subsequently lowering performance [47]. Importantly, when the sample size reaches 10, both GA and PA decrease compared to a sample size of eight. This decrease is attributed to prompt truncation caused by an overload of retrieved logs, which exceeds the context size of the LLM, resulting in incomplete input data.
@@
Figure 4: GA and PA of OpenLogParser using different numbers of selected logs in the prompt.
@@
Using Jaccard similarity for log selection and LLM self-reflection enhances the parsing result, although they come with added overhead. Retrieving more logs within the prompt does not necessarily increase effectiveness; in fact, the optimal number of logs enables OpenLogParser to reach peak accuracy is three.
RQ4: What is the effectiveness of OpenLogParser with different LLMs?
@@
Motivation. Unlike previous parsers [23, 20, 44] that are based on commercial LLMs, OpenLogParser employs open-source LLM to mitigate privacy concerns and monetary costs. Different LLMs exhibit varying capabilities due to their distinct architectures and pre-training data. In this RQ, we evaluate the performance of OpenLogParser across various open-source LLMs.
@@
TABLE IV: Parsing performance of OpenLogParser using different LLMs.
@@ VI Threats to Validity
External validity. Data leakage is a potential risk of LLM-based log parsers [31, 20]. Although OpenLogParser does not involve using labeled logs for fine-tuning or in-context learning, there is a possibility that the LLM might have been pre-trained on publicly available log data. Our evaluation dataset with ground-truth templates was released on August 2023 [19] and Llama3-8B training knowledge cutoff from March 2023 [4], so the leakage risk should be minimal. The log format may also affect our result, but the datasets used are large and cover logs from various systems in different formats. Future studies are needed to evaluate OpenLogParser on logs from other systems.
Internal validity. OpenLogParser employs Llama3-8B as its base model due to its promising results in many tasks and the relatively small size [4]. We also compared the results across various open-source LLMs and found differences. Future research is needed to evaluate LLM-based parsers’ performance when more advanced LLMs are released in the future. The effectiveness of OpenLogParser could be influenced by specific parameter settings (e.g., the number of logs selected for prompting). Our evaluations showed that these settings have an impact on the parsing results and discussed the optimal settings. Future studies are needed to evaluate the settings on other datasets.
Construct validity. To mitigate the effects of randomness in evaluating OpenLogParser, the generation temperature of the model is set to zero. This adjustment ensures that experiments conducted under the same conditions are repeatable and that the results are stable.
@@ VII Conclusion
In this paper, we introduced OpenLogParser, an unsupervised log parsing technique utilizing open-source LLMs to effectively address the limitations of existing LLM-based and syntax-based parsers. OpenLogParser first groups logs that share a syntactic similarity in the static text but vary in the dynamic variable, using a fixed-depth grouping tree. It then parses logs in these groups with three components: i) retrieval augmented generation using similarity scoring: identifies diverse logs within each group based on Jaccard similarity, aiding the LLM in differentiating static text from dynamic variables; ii) self-reflection: iteratively queries LLMs to refine log templates and enhance parsing accuracy; and iii) log template memory: store parsed templates to minimize LLM queries, thereby boosting parsing efficiency. Our comprehensive evaluations on LogHub-2.0, a public large-scale log dataset, demonstrate that OpenLogParser achieves an average GA of 0.8720 and an average PA of 0.8538, outperforming state-of-the-art parsers (i.e., ILIAC [20] and LLMParser [31]) by 5% and 25%, respectively. OpenLogParser parses logs from all 14 systems (50 million logs) in a total of 5.94 hours, which is 2.75 and 40 times faster than other LLM-based parsers This marks a substantial advancement over traditional semantic-based and LLM-based parsers in an unsupervised way, confirming the robustness and effectiveness of our approach. Additionally, OpenLogParser addresses the privacy and cost concerns associated with commercial LLMs, making it a highly efficient and secure solution for practical log parsing needs.

### v3

@@ V Evaluation
During retrieval augmented log parsing, varying the number of selected logs affects the performance of LibreLog. Retrieving three logs into the prompt yields the highest effectiveness. Fig 4 shows the LibreLog performance with variations in the number of logs from a group retrieved into the prompts. LibreLog maintains high accuracy across various sample sizes, with optimal performance achieved when the sample size is set to three, reaching peak values in both GA and PA. Notably, when the sample size is reduced to one, GA and PA drop to 0.80 and 0.70, respectively, representing a decline of 8.26% and 18% compared to a sample size of three. This reduction highlights the challenges LLM faces in parsing logs accurately without sufficient comparative data, such as multiple log comparisons or labeled logs. As the sample size increases from one to two, both GA and PA show significant improvements, peaking when the sample size reaches three. However, further increases in sample size from three to eight result in slight decreases in GA and PA, stabilizing around 0.865 and 0.835, respectively. This suggests that an excess of log samples may introduce noise, subsequently lowering performance [47]. Importantly, when the sample size reaches 10, both GA and PA decrease compared to a sample size of eight. This decrease is attributed to prompt truncation caused by an overload of retrieved logs, which exceeds the context size of the LLM, resulting in incomplete input data.
@@
Figure 4: GA and PA of LibreLog using different numbers of selected logs in the prompt.
@@
Using Jaccard similarity for log selection and LLM self-reflection enhances the parsing result, although they come with added overhead. Retrieving more logs within the prompt does not necessarily increase effectiveness; in fact, the optimal number of logs enables LibreLog to reach peak accuracy is three.
RQ4: What is the effectiveness of LibreLog with different LLMs?
@@
Motivation. Unlike previous parsers [23, 20, 44] that are based on commercial LLMs, LibreLog employs open-source LLM to mitigate privacy concerns and monetary costs. Different LLMs exhibit varying capabilities due to their distinct architectures and pre-training data. In this RQ, we evaluate the performance of LibreLog across various open-source LLMs.
@@
TABLE IV: Parsing performance of LibreLog using different LLMs.
@@ VI Threats to Validity
External validity. Data leakage is a potential risk of LLM-based log parsers [31, 20]. Although LibreLog does not involve using labeled logs for fine-tuning or in-context learning, there is a possibility that the LLM might have been pre-trained on publicly available log data. Our evaluation dataset with ground-truth templates was released on August 2023 [19] and Llama3-8B training knowledge cutoff from March 2023 [4], so the leakage risk should be minimal. The log format may also affect our result, but the datasets used are large and cover logs from various systems in different formats. Future studies are needed to evaluate LibreLog on logs from other systems.
Internal validity. LibreLog employs Llama3-8B as its base model due to its promising results in many tasks and the relatively small size [4]. We also compared the results across various open-source LLMs and found differences. Future research is needed to evaluate LLM-based parsers’ performance when more advanced LLMs are released in the future. The effectiveness of LibreLog could be influenced by specific parameter settings (e.g., the number of logs selected for prompting). Our evaluations showed that these settings have an impact on the parsing results and discussed the optimal settings. Future studies are needed to evaluate the settings on other datasets.
Construct validity. To mitigate the effects of randomness in evaluating LibreLog, the generation temperature of the model is set to zero. This adjustment ensures that experiments conducted under the same conditions are repeatable and that the results are stable.
@@ VII Conclusion
In this paper, we introduced LibreLog, an unsupervised log parsing technique utilizing open-source LLMs to effectively address the limitations of existing LLM-based and syntax-based parsers. LibreLog first groups logs that share a syntactic similarity in the static text but vary in the dynamic variable, using a fixed-depth grouping tree. It then parses logs in these groups with three components: i) retrieval augmented generation using similarity scoring: identifies diverse logs within each group based on Jaccard similarity, aiding the LLM in differentiating static text from dynamic variables; ii) self-reflection: iteratively queries LLMs to refine log templates and enhance parsing accuracy; and iii) log template memory: store parsed templates to minimize LLM queries, thereby boosting parsing efficiency. Our comprehensive evaluations on LogHub-2.0, a public large-scale log dataset, demonstrate that LibreLog achieves an average GA of 0.8720 and an average PA of 0.8538, outperforming state-of-the-art parsers (i.e., ILIAC [20] and LLMParser [31]) by 5% and 25%, respectively. LibreLog parses logs from all 14 systems (50 million logs) in a total of 5.94 hours, which is 2.75 and 40 times faster than other LLM-based parsers This marks a substantial advancement over traditional semantic-based and LLM-based parsers in an unsupervised way, confirming the robustness and effectiveness of our approach. Additionally, LibreLog addresses the privacy and cost concerns associated with commercial LLMs, making it a highly efficient and secure solution for practical log parsing needs.